OpenStudy (pokemon23):

The length of a rectangle is 3 times its width. The perimeter of rectangle is 72 meters. Find the dimensions of the rectangle.

OpenStudy (across):

l = 3w, P = 2(l+w)=72, 2(l+w)=72, 2(3w+w)=72, 2(4w)=72, 8w=72, w=9. *** l=3w, l=3(9), l=27. ***